# Tree Ensembles for Contextual Bandits

### v2

@@ 4.3.2 Experimental Results
Fig. 2 shows the results of the TEUCB and TETS methods along with the baselines. The road network and the agents’ traversals are presented in Fig. 3. The frequently an edge has been traversed is indicated by the red saturation, where more traversals correspond to a higher saturation. It is worth noting that LinUCB, LinTS, and the TreeBootstrap require excessive exploration as their models are edge-specific. On the contrary, TETS, TEUCB, NeuralTS, and NeuralUCB train one model with the possibility of generalizing the expected travel time predictions over different edges. We observe that using both XGBoost and random forest, TEUCB and TETS significantly outperform the other methods. In this setting, if an agent can generalize well what it has learned from one arm’s reward distribution to the other arms, it then can effectively avoid over-exploration, as indicated by both the regret plots in Fig. 2 and the agent’s specific path selections in Fig. 3. Comparing the two different tree ensemble methods, XGBoost seems to perform better on the navigation task compared to random forest in the TEUCB and TETS frameworks.
@@
All the agents were run on a single desktop CPU, apart from NeuralUCB and NeuralTS, which were run on an NVIDIA A40 GPU. NeuralTS and NeuralUCB tend to require large networks and substantial hyper-parameter tuning to achieve deliberate exploration, which can make them time-consuming while running and searching for good parameter values. In terms of run-time, the experiments took about an hour on the GPU for a neural agent. It is also noticeable that the neural agents appear to be sensitive to the weight initialization of the networks, which leads to higher variance (see Fig. 2). In contrast, TETS and TEUCB achieve solid results for a large range of parameter settings, yield smaller regret, and tend to run at comparable speed on a single CPU. The experiments with TEUCB and TETS took about 1.5 hours with XGBoost, and 5 hours using random forest as the tree ensemble method. The run-time of TreeBootstrap also tends to depend heavily on the particular tree model used. With a single decision tree per arm, the experiments took 0.5 hour, which is about the same as LinUCB and LinTS. Using tree ensembles, they took about 4 and 20 hours with XGBoost and random forests, respectively.
We note that TreeBootstrap with tree ensembles may have lower runtimes. Elmachtoub et al. (2017) propose a trick to speed up this method with a computational shortcut called TreeHeuristic. However, as acknowledged, its effectiveness is affected by some shortcomings including the inability to generalize over different but related arms through contextual information.
@@ 5 Conclusion
We developed a novel framework for contextual multi-armed bandits using tree ensembles. Within this framework, we incorporated the two commonly used methods for handling the exploration-exploitation dilemma: UCB and Thompson Sampling. Furthermore, we extended the framework to handle combinatorial contextual bandits, enabling more complex action selection at each time step.
To demonstrate the effectiveness of the framework, we conducted extensive experiments on benchmark datasets using the XGBoost and random forest algorithms. Additionally, we utilized it for navigation over stochastic real-world road networks.

### v3

@@ 4.3.2 Experimental Results
Fig. 2 shows the results of the TEUCB and TETS methods along with the baselines. The road network and the agents’ traversals are presented in Fig. 3 (in the Appendix). The frequency by which an edge has been traversed is indicated by the red saturation, where more traversals correspond to a higher saturation. It is worth noting that LinUCB, LinTS, and the TreeBootstrap require excessive exploration as their models are edge-specific. In contrast, TETS, TEUCB, NeuralTS, and NeuralUCB train one model with the possibility of generalizing the expected travel time predictions over different edges. We observe that using both XGBoost and random forest, TEUCB and TETS significantly outperform the other methods. In this setting, if an agent can generalize well what it has learned from one arm’s reward distribution to the other arms, it can then effectively avoid over-exploration, as indicated by both the regret plots in Fig. 2 and the agent’s specific path selections in Fig. 3. Comparing the two different tree ensemble methods, XGBoost seems to perform better on the navigation task compared to random forests in the TEUCB and TETS frameworks.
@@
All the agents were run on a single desktop CPU, apart from NeuralUCB and NeuralTS, which were run on an NVIDIA A40 GPU. NeuralTS and NeuralUCB tend to require highly parameterized networks and substantial hyper-parameter tuning to achieve deliberate exploration, which can make them time-consuming while running and searching for good parameter values. In terms of runtime, the experiments took about an hour on the GPU for a neural agent. It is also noticeable that the neural agents appear to be sensitive to the weight initialization of the networks, which leads to higher variance (see Fig. 2). In contrast, TETS and TEUCB achieve solid results for a large range of parameter settings, yield lower regret, and tend to run at comparable speed on a single CPU (instead of GPU). The experiments with TEUCB and TETS took about 1.5 hours with XGBoost, and 5 hours using random forests as the tree ensemble methods. The runtime of TreeBootstrap also tends to depend heavily on the particular tree model used. With a single decision tree per arm, the experiments took 0.5 hours, which is about the same as LinUCB and LinTS. Using tree ensembles, they took about 4 and 20 hours with XGBoost and random forests, respectively.
@@ 5 Conclusion
We developed a novel framework for contextual multi-armed bandits using tree ensembles. Within this framework, we adapted the two commonly used methods for handling the exploration-exploitation dilemma: UCB and Thompson Sampling. Furthermore, we extended the framework to handle combinatorial contextual bandits, enabling more complex action selection at each time step. To demonstrate the effectiveness of the framework, we conducted experiments on benchmark datasets using the XGBoost and random forest tree ensemble methods. Additionally, we employed it for navigation over stochastic real-world road networks, modeled as a combinatorial contextual bandit problem.
Across all problem instances, the introduced tree ensemble-based methods, TEUCB and TETS, consistently demonstrated effectiveness in minimizing regret while requiring relatively low computational resources. In many cases, these methods significantly outperformed neural network-based methods that are often considered state-of-the-art for contextual bandits, in terms of both accuracy and computational cost. Hence, this work indicates that using tree ensembles offers several advantages over other machine-learning models for these problems. Furthermore, compared to other tree-based methods, TEUCB and TETS exhibited similar or better capabilities in minimizing regret while maintaining computational efficiency, even for problems with large numbers of arms. These results indicate that using a single tree ensemble to predict the reward distributions of all arms, particularly as implemented in TEUCB and TETS, facilitates effective generalization between arms based on context, enabling more efficient learning with fewer samples.
The study primarily focuses on the practical applicability of tree ensembles for bandits in various settings. Therefore, we did not provide a theoretical analysis of the respective regrets. For future work, exploring theoretical regret bounds for the proposed tree ensemble methods could offer deeper insights and a better understanding of the methods presented in this work.
@@ Acknowledgments
Appendix A Appendix
A.1 Nomenclature
Table 4: Table of notations used in this paper.
Notation Description
$a$ Arm/action
$a^{*}$ Optimal arm
$a_{t}$ Arm selected at time step $t$
$\mathcal{A}$ Action space, set of all arms
$c_{n}$ Number of training samples assigned to a leaf in regression tree $n$
$f$ Tree ensemble regressor consisting of $N$ regression trees fitted in unison
$f_{n}$ $n$th individual tree of the ensemble
$m_{t,a}$ Number of times arm $a$ has been selected up to time step $t$
$\mathcal{N}(\mu,\sigma^{2})$ Normal distribution with mean $\mu$ and variance $\sigma^{2}$
$o_{n}$ Output value associated with a leaf in regression tree $n$
$q$ Function of expected reward
$r$ Observed reward
$\tilde{r}$ Estimated reward
$R$ Regret
$s_{n}^{2}$ Sample variance of the output value associated with a leaf in regression tree $n$
$\mathcal{S}$ Super arm, a set of arms
$\mathcal{S}^{*}$ Optimal super arm
$\mathcal{S}_{t}$ super arm selected at time step $t$
$t$ time step
$U$ Upper confidence bound
$\mathbf{x}$ Context/feature vector
$\mu$ True mean
$\tilde{\mu}$ Estimated mean
$\nu$ Exploration factor
$\sigma^{2}$ True variance
$\tilde{\sigma}^{2}$ Estimated variance
$\{x_{i}\}_{i=1}^{I}$ List of $I$ entries
A.2 Further Elaboration on the Independence Assumption
Given a tree ensemble $f=\{f_{1},f_{2},\dots,f_{N}\}$, the prediction for a context vector $\mathbf{x}$ is the sum of individual tree outputs:
$p(\mathbf{x})=\sum_{i=1}^{N}o_{i}(\mathbf{x}),$ (7)
where each $o_{i}(\mathbf{x})$ is a random variable representing the output of tree $i$ when the input is $\mathbf{x}$. Assuming the output of each tree $o_{i}(\mathbf{x})$ has a mean $\mu_{i}$ and a variance $\frac{\sigma_{i}^{2}}{c_{i}}$ (where $c_{i}$ is the number of training samples assigned to the leaf of tree $i$ that $\mathbf{x}$ ends up in), the variance of the ensemble’s prediction is:
$\text{Var}(p(\mathbf{x}))=\text{Var}\left(\sum_{i=1}^{N}o_{i}(\mathbf{x})\right).$ (8)
Under the independence assumption between the trees, the variance of the sum is simply the sum of the variances of the individual trees:
$\text{Var}(p(\mathbf{x}))=\sum_{i=1}^{N}\text{Var}(o_{i}(\mathbf{x}))=\sum_{i=1}^{N}\frac{\sigma_{i}^{2}}{c_{i}}.$ (9)
This expression allows us to estimate the total uncertainty in the ensemble’s prediction by summing up the individual uncertainties of each tree.
In reality, especially in boosting algorithms, there might be some correlation between the trees. Let us denote the correlation between the outputs of two trees $i$ and $j$ as $\rho_{ij}$. The variance of the sum of tree outputs in the presence of correlations becomes:
$\text{Var}(p(\mathbf{x}))=\sum_{i=1}^{N}\frac{\sigma_{i}^{2}}{c_{i}}+2\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\rho_{ij}\sqrt{\frac{\sigma_{i}^{2}}{c_{i}}\cdot\frac{\sigma_{j}^{2}}{c_{j}}}.$ (10)
Here, $\rho_{ij}$ represents the correlation coefficient between the outputs of trees $i$ and $j$. If the correlations $\rho_{ij}$ are positive, they increase the total variance of the ensemble’s prediction. However, if we assume independence ($\rho_{ij}=0$), the second term vanishes, simplifying the variance to just the sum of individual variances, i.e.,
$\text{Var}(p(\mathbf{x}))\approx\sum_{i=1}^{N}\frac{\sigma_{i}^{2}}{c_{i}}.$ (11)
In the context of bagging-based ensembles like random forests, where trees are trained on different subsets of data and with random feature selection, the correlations $\rho_{ij}$ are typically small, making this approximation reasonable. For boosting-based methods, which sequentially fit trees, the correlations might be stronger, implying that our approximation could be less accurate.
However, we argue that this approximation often yields a lower bound for the actual variance, where this lower bound provides useful insights into the ensemble’s uncertainty, offering a simplified way to quantify prediction variability. For this, we argue that negative correlations are less common than positive ones in ensemble learning primarily due to the nature of how individual models are trained and how ensemble methods perform.
In tree ensemble methods, individual models (e.g., decision trees) are trained on subsets of the same data and tend to capture underlying patterns that are common across these subsets. This shared learning process often leads to similar predictions, resulting in positive correlations. In boosting, in particular, trees are added sequentially, with each new tree focused on correcting the errors (residuals) of the previous ensemble. As the boosting process continues, new trees are highly dependent on the earlier trees’ mistakes. Although boosting can sometimes create predictions that ‘pull’ in opposite directions (implying a negative correlation locally), the overall direction of learning is typically aligned to reduce residual error. This alignment contributes to positive correlations because all trees are ultimately working together to approximate the same target function. Therefore, we can assume that in boosting, the overall correlation is positive. On the other hand, as mentioned, bagging methods such as random forests use randomness (bootstrapping data samples and feature selection) to reduce the correlation between individual trees (i.e., the correlations can be discarded).
In summary, we expect the independence assumption between the trees (i.e., $\rho_{ij}=0$) to hold reasonably well for bagging methods. On the other hand, this assumption usually leads to an underestimate of the total variance for boosting methods. This means that our calculated variance: $\sum_{i=1}^{N}\frac{\sigma_{i}^{2}}{c_{i}}$ serves as a lower bound on the true variance. Interestingly, some bandit algorithms, such as UCB, are known to suffer from over-exploration [Russo and Roy, 2014], and thus underestimating the variance could even be helpful to mitigate this issue and lead to more efficient learning.
A.3 Calculating Leaf Values
The code blocks presented in this section outline the procedure by which the leaf values $o_{n,l}$, $s^{2}_{n,l}$ and $c_{n,l}$ for a leaf $l$ in tree $n$ are obtained for the implementation of TEUCB and TETS with XGBoost and random forests, respectively.
Algorithm 2 Set Leaf Values - XGBoost
1:  Input: Tree ensemble regressor $f$, number of trees in ensemble $N$, base predictor value $b$, learning rate $\eta$, set of $D$ training data points $\left\{(\mathbf{x}_{i},r_{i})\right\}_{i=1}^{D}$.
2:  $o_{n,l},s_{n,l},c_{n,l}=0\quad\forall\text{ trees }n,\text{ leaves }l$
3:  for $i=1$ to $D$ do
4:     for $n=1$ to $N$ do
5:        Check which leaf $l$ that the context $\mathbf{x}_{i}$ is assigned to in tree $n$
6:        $c_{n,l}=c_{n,l}+1$
7:        Get prediction on $\mathbf{x}_{i}$ from previous trees: $p_{n-1}=b+\sum_{j=1}^{n-1}f_{j}(\mathbf{x}_{i})$
8:        $o_{n,l,c_{n,l}}=\eta(r_{i}-p_{n-1})$
9:     end for
10:  end for
11:  for $n=1$ to $N$ do
12:     for leaves $l$ in tree $n$ do
13:        $o_{n,l}=\frac{1}{c_{n,l}}\sum_{i=1}^{c_{n,l}}o_{n,l,i}$
14:        $s_{n,l}^{2}=\frac{1}{c_{n,l}-1}\sum_{i=1}^{c_{n,l}}(o_{n,l,i}-o_{n,l})^{2}$
15:     end for
16:  end for
Algorithm 3 Set Leaf Values - Random Forest
1:  Input: Tree ensemble regressor $f$, number of trees in ensemble $N$, set of $D$ training data points $\left\{(\mathbf{x}_{i},r_{i})\right\}_{i=1}^{D}$.
2:  $o_{n,l},s_{n,l},c_{n,l}=0\quad\forall\text{ trees }n,\text{ leaves }l$
3:  for $i=1$ to $D$ do
4:     for $n=1$ to $N$ do
5:        Check which leaf $l$ that the context $\mathbf{x}_{i}$ is assigned to in tree $n$
6:        $c_{n,l}=c_{n,l}+1$
7:        $o_{n,l,c_{n,l}}=f_{j}(\mathbf{x}_{i})$
8:     end for
9:  end for
10:  for $n=1$ to $N$ do
11:     for leaves $l$ in tree $n$ do
12:        $o_{n,l}=\frac{1}{c_{n,l}}\sum_{i=1}^{c_{n,l}}o_{n,l,i}$
13:        $s_{n,l}^{2}=\frac{1}{c_{n,l}-1}\sum_{i=1}^{c_{n,l}}(o_{n,l,i}-o_{n,l})$
14:     end for
15:  end for
A.4 Path Selections
In this section, we present maps of the routes selected by the respective agents throughout one run of the navigation task.
Figure 3: The road network of Luxembourg shows the trajectories of different agents for the experiment on problem instance 1, where the corresponding cumulative regret is presented in Fig. 2(a). The plots show all the paths selected by the agents during a full run of the experiment, where a higher level of opacity indicates that a road segment was more frequently part of a traveled path.
A.5 Sensitivity Analysis
Here, we investigate the robustness of TEUCB and TETS to changes in a number of key (hyper)parameters. To do so, we run the same experiments as presented in Section 4.2, but varying one parameter at a time. In addition, we study how robust the methods are to delays in reward feedback, and compare the performance of TEUCB and TETS to TreeBootstrap.
A.5.1 Delayed Feedback
As done by Zhang et al. [2021], in the experiments with delayed feedback presented in Fig. 4, the rewards associated with an agent’s actions are fed in batches of varying batch size, rather than one by one instantly after the action is taken. This setting appears to occur naturally in several real-world applications of multi-armed bandits, as discussed by Chapelle and Li [2011]. Their findings suggest that Thompson Sampling scales better with delay in rewards than UCB methods, which is also supported by the experiments of Zhang et al. [2021]. It is not clear from our experiments whether this is the case for TEUCB, TETS, and TreeBootstrap (which, as discussed in Section 1.1, can be viewed as a Thompson Sampling method). Comparing with the results of Zhang et al. [2021] for NeuralUCB and NeuralTS on the same datasets, however, it appears that the tree-based methods we study are more resilient to delays in the rewards.
Figure 4: Comparison of TEUCB, TETS, and TreeBootstrap on the mushroom and the adult datasets with different levels of delays. Cumulative regret for a single experiment is calculated over 10,000 time steps and repeated 10 times. The results display the average cumulative regret plus/minus the standard deviation for each respective agent. Hyperparameters are selected as in Section 4.2. All agents are evaluated on the same levels of reward delays, but each agent is shifted slightly horizontally for visualization purposes.
A.5.2 Varied Tree Depth
In Fig. 5, we study the robustness of TEUCB and TETS towards different depths for the regression trees in the ensembles. We observe that a significant change in tree depth may influence the performance of TEUCB and TETS when used with both XGBoost and random forests. We note that random forests tend to benefit from deeper trees, while XGBoost performs better with shallower ones. This is particularly interesting as the depth is only a maximum value for XGBoost, and the algorithm is designed to automatically prune trees where deemed beneficial for the reduction of overfitting. The extent to which this is utilized can be controlled by changing the XGBoost parameters $\gamma$ and $\lambda$, though we use the default values. As a final note on the effect of tree depth, we observe that a depth of 10, which is used in the comparison study in Section 4.2, is suitable.
Figure 5: Comparison of TEUCB and TETS on the mushroom and the adult datasets with different tree depths. Cumulative regret for a single experiment is calculated over 10,000 time steps and repeated 10 times. The results display the average cumulative regret plus/minus the standard deviation for each respective agent. Hyperparameters other than tree depth are selected as in Section 4.2. All agents are evaluated with the same tree depths, but each agent is shifted slightly horizontally for visualization purposes.
A.5.3 Varied Ensemble Size
In Fig. 6 we study how the performance of TEUCB and TETS varies with different numbers of trees making up the tree ensembles used to predict rewards. In this study, we see that initially going from a relatively small ensemble of trees to a larger one can greatly improve the performance, which makes sense. However, after this initial setup, the results stay stable when increasing the number of trees further. XGBoost appears to be particularly sensitive to selecting too few trees. This issue may potentially be mitigated by adjusting the learning rate parameter $\eta$, though we have not investigated this. Further, we note that setting the number of trees to 100, as in the comparison with other bandit methods in Section 4.2, is shown to be a good choice.
Figure 6: Comparison of TEUCB and TETS on the mushroom and the adult datasets with different numbers of trees in the ensembles. Cumulative regret for a single experiment is calculated over 10,000 time steps and repeated 10 times. The results display the average cumulative regret plus/minus the standard deviation for each respective agent. Hyperparameters other than the numbers of trees are selected as in Section 4.2. All agents are evaluated with the same number of trees, but each agent is shifted slightly horizontally for visualization purposes.
A.5.4 Varied Exploration Factor
Finally, in Fig. 7 we investigate the robustness of TEUCB and TETS to different exploration factors used to control the level of exploration. The results demonstrate that the methods are robust to changes in the exploration factor within a wide range of values. Only one observation deviates significantly from the pattern, in which an exploration factor of 100 is used for TEUCB with XGBoost. For both datasets, this value results in a significantly larger average cumulative regret compared to all other evaluated methods and parameter values. For all other values and agents, the results are stable and consistent.
Figure 7: Comparison of TEUCB and TETS on the mushroom and the adult datasets with different exploration factors. Cumulative regret for a single experiment is calculated over 10,000 time steps and repeated 10 times. The results display the average cumulative regret plus/minus the standard deviation for each respective agent. Hyperparameters other than the exploration factor are selected as in Section 4.2. All agents are evaluated on the same exploration factors, but each agent is shifted slightly horizontally for visualization purposes.